in an isosceles right teiangle the length of the hypotenuse is 10 units. how long is each leg?
Answer by brysca(112)

Okay so since the sides are equal to each other, you can say that both sides equal the same value, x.
Now we use the pythagorean theorem and substitute a and b for x.
a^2+b^2=c^2
x^2+ x^2=10^2
2x^2=100
Divide by 2.
x^2=50
Take the square root of both sides.
x=√50
x=5√2
Each side is equal to 5√2.